# ITEACNet: Inverted Teacher-studEnt seArch Conversation Network

### v2

@@ Results
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 Average
IEMOCAP (four-class) CPM-Net 44.05 44.33 42.63 41.60 41.18 41.48 41.21 41.37 42.23
AE 41.13 42.15 42.50 43.14 43.33 43.61 42.60 42.73 42.65
CRA 33.56 32.66 31.97 31.46 30.24 28.76 27.95 27.04 30.46
MMIN 70.80 69.25 67.25 64.93 62.32 59.60 56.93 54.21 63.16
GCNet 64.90 70.13 72.60 73.69 74.83 74.99 74.75 75.05 72.62
ITEACNet 77.13 76.46 76.08 75.67 74.50 73.88 73.17 72.78 74.96
IEMOCAP (four-class) CPM-Net 28.49 28.18 26.92 26.14 25.99 26.24 26.05 26.63 26.83
AE 24.60 24.41 24.55 24.04 23.87 23.05 23.23 22.63 23.80
CRA 16.16 23.69 23.14 22.63 21.94 20.98 19.73 19.36 20.95
MMIN 49.30 47.76 46.61 44.74 42.96 41.43 39.58 38.07 43.81
GCNet 44.39 49.31 52.88 54.24 55.20 54.42 54.52 54.21 52.40
ITEACNet 58.63 58.19 58.08 56.97 56.95 55.86 55.01 54.30 56.75
CMUMOSI CPM-Net 66.51 68.49 66.97 65.78 66.18 65.79 65.28 64.43 66.18
AE 62.36 60.37 59.58 57.57 56.49 53.62 51.56 50.03 56.45
CRA 61.27 58.54 55.78 54.24 50.90 48.16 45.92 41.80 52.08
MMIN 83.71 81.01 78.66 76.27 72.00 68.20 65.19 62.11 73.39
GCNet 68.79 68.17 70.47 70.63 68.59 69.47 67.88 69.55 69.19
ITEACNet 85.38 83.67 81.98 78.74 77.16 71.61 70.63 66.99 77.02
CMUMOSEI CPM-Net 74.17 74.58 74.51 74.06 72.83 71.32 69.67 68.26 72.43
AE 78.16 77.20 76.41 75.89 74.80 73.80 72.64 71.43 75.04
CRA 80.65 78.50 75.80 73.73 71.46 68.88 66.19 63.32 72.32
MMIN 85.24 83.02 80.30 78.08 74.90 72.08 67.86 65.85 75.91
GCNet 78.69 80.48 80.99 80.95 80.85 80.93 80.32 79.79 80.38
ITEACNet 87.08 86.15 85.38 84.43 83.17 81.79 80.92 79.90 83.60
Inverted Teacher-Student Traning Performance
Dealing with incomplete data poses a significantly greater challenge than handling complete data. However, current methods do not equip models for handling incomplete data with more stronger abilities to follow the performance of models trained with complete data. Therefore, we propose the Inverted Teacher-Student traning framework, which employs a student model with complex structures to handle incomplete data while emulating the performance of a teacher model with simpler structures trained with complete data. To comprehensively analyze the performance of ITS, we compare different training frameworks under different settings. The results are presented in Table 2, where the ”GCNet” training framework involves the teacher model receiving complete and incomplete data during the training phase, with gradients accumulated from both types of data and updated together at each step. The ”TS” training framework represents the vanilla Teacher-Student framework, where the teacher model is more complex than the student model. ”TF” is an abbreviation for Transformer. It’s worth noting that, in the TS training framework, the teacher model is equipped with four encoder layers, while the student model is equipped with three.
Firstly, the performance of the Transformer under GCNet training framework is not satisfactory. It not only fails to achieve outstanding performance with complete data but also exhibits poor capability in handling incomplete data. The IST training framework brings about a noticeable improvement in performance. The student model, guided by a simpler teacher model, not only performs better with complete data, surpassing the baselines across all four tasks, but also demonstrates significant performance improvements in handling incomplete data. In transitioning from complete data to a severe missing scenario with a 70% data loss, the ITS training framework exhibits less performance degradation. Compared to the GCNet training framework, ITS incurs a smaller loss of 2.88% in the four-class IEMOCAP task, 5.24% in the six-class IEMOCAP task, and 1.28% in the CMUMOSI task. While it shows a slight disadvantage in the CMUMOSEI task, the difference is negligible due to the similar performance under various missing rates for both frameworks.
Furthermore, we propose the efficient augmentation of the student model’s complexity using Neural Architecture Search (NAS), aiming to enhance its capability in handling incomplete data. While the introduction of NAS doesn’t necessarily result in superior performance on complete data, it significantly enhances the model’s capability to handle incomplete data. The student model enhanced through NAS achieved the smallest performance decline across all four tasks, with reductions of 3.63% in the four-class IEMOCAP, 3.08% in the six-class IEMOCAP, 15.97% in CMUMOSI, and 6.50% in CMUMOSEI. This substantial surpasses the previous settings. These results effectively demonstrate that by strengthening the computational capacity of the student model in ITS, it can handle incomplete data more effectively, even when trained with the same teacher.
Table 5: Model Params
Model Module Parameters
CPM-Net - 0.57M
AE - 1.29M
CRA - 2.11M
MMIN - 3.87M
GCNet - 34.14M
ITEACNet Teacher 2.19M
Student 3.88M
ITEACNet Performance
Independent Model Evaluation Performance
To demonstrate the outstanding performance of ITEACNet in CER tasks, we maintained the same evaluation method as previous works in Table 3, enabling a fair comparison with the baselines.
Firstly, the Average metric reflect the overall performance of the model under different missing scenarios. ITEACNet surpasses the current state-of-the-art (sota) work, GCNet, by 0.52%, 1.38%, and 0.17% in the four-class IEMOCAP, six-class IEMOCAP, and CMUMOSI tasks, respectively, achieving the new sota performance. Although ITEACNet’s performance in the CMUMOSEI task is 0.19% lower than GCNet, considering the carefully design of GCNet’s network structure, this performance remains highly competitive.
Furthermore, observing ITEACNet’s performance under different missing rates, we find that despite GCNet utilizing complex network structures to additionally process speaker information in conversation, there is still no significant difference between our method and GCNet across all four tasks. Even in severe missing scenarios, such as at a 70% missing rate, our method demonstrates more pronounced advantages in the four-class IEMOCAP and six-class IEMOCAP tasks, achieving performance improvements of 2.46% and 2.48%, respectively. These results collectively demonstrate the excellent robustness of ITEACNet.
Unified Model Evaluation Performance
To more accurately demonstrate the model’s robustness in real-world scenarios, we further employ the UME method to evaluate their performance.. The experimental results are presented in Table 4. As previous works did not provide experimental results for this evaluation method, all results presented here are reproduced by us. The parameter quantities for each model during the reproduction process are listed in Table 5.
Firstly, under the Average metric, ITEACNet exhibits a significant performance advantage across all four tasks. Compared to the state-of-the-art methods, our approach achieves performance improvements of 2.34%, 4.35%, 7.83%, and 3.22% in the four-class IEMOCAP, six-class IEMOCAP, CMUMOSI, and CMUMOSEI tasks, respectively.
Furthermore, observing ITEACNet’s performance under different missing rates reveals that despite the student model being exposed only to incomplete data during training, it still maintains excellent handling capabilities for complete data. This is attributed to the knowledge transferred from the teacher model, which processes complete data. In contrast, most baseline methods, except for the distillation-based MMIN, lack the ability to handle complete data, resulting in a substantial decrease in Average metric. Simultaneously, the student model in ITEACNet, utilizing an equivalent parameter quantity, achieves significantly superior performance compared to MMIN. This further underscores the superiority of our method.
Conclusion
@@
In this paper, we propose a novel framework, ITEACNet, for incomplete multimodal learning in CER. ITEACNet has a new ECCE module to capture emotional context changes, and experimental results verify the significance of context changes in CER. It also includes a novel ITS training framework, allowing a complex student model to leverage incomplete data, following the performance of a simple teacher model trained on complete data. This design provides the student model with sufficient capability to handle incomplete data. We employ NAS to enhance the complexity of the student model, demonstrating the importance of a complex student model in incomplete learning. Finally, using an evaluation method closer to real-world scenarios, we show that previous methods lose the ability to handle complete data when dealing with incomplete data, while highlighting the importance of the ITS framework in incomplete learning.

### v3

@@ 5.4.2 Unified Model Evaluation Performance
[…]
IEMOCAP (four-class) AE[54] 67.60 66.23 64.09 62.28 59.91 56.97 53.73 51.41 60.28
CRA[41] 61.34 59.12 56.59 53.90 51.12 48.10 43.98 41.51 51.96
MMIN[11] 74.77 72.74 70.34 67.59 64.98 61.81 58.30 56.05 65.82
GCNet[13] 71.67 73.53 74.90 75.96 75.72 76.20 75.48 74.80 74.78
ITEACH-Net 77.47 76.86 76.78 75.86 75.26 74.33 73.73 71.79 75.26
IEMOCAP (six-class) AE[54] 48.08 46.38 44.71 43.20 41.08 39.07 36.37 34.28 41.64
CRA[41] 36.01 34.95 34.07 32.60 31.17 29.68 26.51 25.24 31.28
MMIN[11] 57.16 54.81 53.05 50.15 47.85 44.79 41.39 38.90 48.51
GCNet[13] 51.88 53.88 55.34 55.53 55.91 56.22 56.16 55.27 55.03
ITEACH-Net 59.09 58.72 58.13 57.53 57.18 56.39 55.53 55.06 57.21
CMUMOSI AE[54] 85.66 81.72 78.29 74.14 69.46 62.81 58.13 54.23 70.55
CRA[41] 85.00 80.90 77.01 72.03 67.46 61.86 56.75 52.24 69.16
MMIN[11] 84.62 79.84 76.32 71.39 68.05 61.95 56.40 52.93 68.94
GCNet[13] 68.78 70.91 70.72 70.11 69.45 69.72 68.28 68.16 69.52
ITEACH-Net 84.40 82.39 80.54 78.86 75.62 74.30 69.92 65.77 76.48
CMUMOSEI AE[54] 85.19 83.14 81.50 79.33 77.00 74.43 71.52 70.19 77.79
CRA[41] 85.73 83.18 80.61 77.51 74.69 71.94 67.84 65.49 75.88
MMIN[11] 85.21 82.83 80.19 78.09 74.53 71.53 67.90 66.08 75.79
GCNet[13] 82.81 82.67 82.72 82.31 82.03 81.81 81.12 79.99 81.93
ITEACH-Net 86.91 85.78 84.86 84.16 82.77 81.60 79.80 78.91 83.10
Progressive Strategy: Compared to handling incomplete data, classifying tasks with complete data is evidently simpler. Hence, we aim to familiarize the model with complete data modeling initially, gradually introducing it to incomplete data in the hope that it can leverage previously acquired knowledge for better handling of incomplete data. Specifically, we start with a 0% missing rate and increase it of the training data by 10% every 10 epochs, up to a maximum of 70%. This progressive strategy allows the model to adapt to incomplete data more smoothly.
As depicted in Table 5, ITEACH-Net consistently achieved optimal results on the Average metrics across all four tasks, with a particularly distinct performance advantage in all tasks except the four-class IEMOCAP. Although GCNet outperforms ITEACH-Net in the four-class IEMOCAP, six-class IEMOCAP and CMU-MOSEI tasks under severe data missing scenarios, it fails to maintain performance at lower missing rates, highlighting a drawback of the previous training frameworks. ITEACH-Net effectively mitigates this issue.
Furthermore, examining the performance of different methods across various missing rates, baselines under the Random training strategy show a balanced performance across different data missing rate scenarios but rarely achieve standout results. With the Progressive training strategy, baselines tend to exhibit a biased proficiency, performing well at a certain missing rate but declining as the missing rate changes, whether it increases or decreases. This indicates that the model is unable to learn how to process data at different missing rates simultaneously. However, regardless of the training strategy employed, ITEACH-Net consistently demonstrates a balanced performance, exhibiting exceptional capability in processing both incomplete and complete data.
5.5 Visualization of Embedding Space
To qualitatively analyze the robustness of ITEACH-Net in the face of dynamic data missing rates, we visualized the latent representations under different missing scenarios on the four-class IEMOCAP test set. Figure 9 presents the t-SNE [55] visualization results. Subfigure 9(a) represents the encoding results of the Teacher model using complete data, while subfigures 9(b) to 9(i) depict the encoding results of the Student model under various data missing rates. It is observable that, in dynamic missing scenarios, the Student model still effectively emulates the knowledge of the Teacher during the encoding complete data. Moreover, as the data missing rate increases, although the variance of the same class data representation gradually enlarges, the overall distribution pattern remains unchanged, and there is still a clear distance between different class data representations. The stable data embedding space distribution demonstrates the robustness of ITEACH-Net in handling different missing scenarios.
(a) Teacher
(b) Student-0.0
(c) Student-0.1
(d) Student-0.2
(e) Student-0.3
(f) Student-0.4
(g) Student-0.5
(h) Student-0.6
(i) Student-0.7
Figure 9: t-SNE visualization results on the four-class IEMOCAP test set with the increasing data missing rates. Student-0.1 represents the results achieved by the Student model when the data missing rate is 10%. We use red, blue, yellow, and green to represent happiness, sadness, neutral, and anger, respectively.
5.6 Visualization of Confusion Matrix
To delve into the processing results of ITEACH-Net on imbalanced category data under different missing rates, we further visualized the confusion matrix of the model’s prediction results in the Four-class IEMOCAP. The results are shown in Figure 10. It can be seen that the prediction accuracy for the Sad, Neutral, and Angry categories decreases with the increase in data missing rates. Among them, the Neutral category has the largest number of samples, totaling 1708, while the Angry category has the least, with only 1103 samples. In contrast, the Happy category, which has a larger number of samples, maintains a relatively stable prediction accuracy under different missing rates. Based on these results, it can be concluded that ITEACH-Net performs stably in the relatively simple task of Happy emotion detection in dynamic missing scenarios, and does not suffer from model collapse due to data imbalance in other emotion detection tasks. Similar prediction distributions under various missing rates also demonstrate the stability of ITEACH-Net in handling different missing scenarios.
(a) Student-0.0
(b) Student-0.2
(c) Student-0.4
(d) Student-0.7
Figure 10: Visualization results of confusion matrices on the four-class IEMOCAP with the increasing data missing rates.
6 Conclusion
@@
In this paper, we propose a novel framework, ITEACH-Net, for incomplete multimodal learning in ERC. ITEACH-Net has a new ECCE module to capture emotional context changes, and experimental results verify the significance of this unique pattern in ERC. It also includes a novel ITS training framework, allowing a complex student model to leverage incomplete data, following the performance of a simple teacher model trained on complete data. This design provides the student model with sufficient capability to handle incomplete data, and experimental results demonstrate the importance of a complex student model in incomplete learning. Finally, using an evaluation method, we demonstrate that previous approaches struggle to effectively handle varying data missing rate scenarios when dealing with incomplete data. Concurrently, experimental results showcase the superiority of ITEACH-Net in incomplete learning.
In the future, we aim to expand ITEACH-Net to focus on real-time analysis scenarios, considering how to determine the emotional context patterns of the current time node based on data from previous time nodes. Additionally, beyond emotion recognition in conversation, we will also utilize ITEACH-Net to address the issue of modality missing in other types of conversation understanding tasks.
7 Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) ( No.62276259, No.62201572, No.U21B2010, No.62271083, No.62306316)
Appendix A Embedding Space
We further provide t-SNE visualization results on other three tasks. Stable embedding distributions at different missing rates demonstrate the model’s stability and robustness when dealing with incomplete data.
(a) Teacher
(b) Student-0.0
(c) Student-0.1
(d) Student-0.2
(e) Student-0.3
(f) Student-0.4
(g) Student-0.5
(h) Student-0.6
(i) Student-0.7
Figure 11: t-SNE visualization results on the Six-class IEMOCAP test set with the increasing data missing rates. Student-0.1 represents the results achieved by the Student model when the data missing rate is 10%. We use red, blue, yellow, green, purple and pink to represent happiness, sadness, neutral, anger, excited and frustrated respectively.
(a) Teacher
(b) Student-0.0
(c) Student-0.1
(d) Student-0.2
(e) Student-0.3
(f) Student-0.4
(g) Student-0.5
(h) Student-0.6
(i) Student-0.7
Figure 12: t-SNE visualization results on the CMU-MOSEI test set with the increasing data missing rates. We use red and blue to represent positive and negtive respectively.
(a) Teacher
(b) Student-0.0
(c) Student-0.1
(d) Student-0.2
(e) Student-0.3
(f) Student-0.4
(g) Student-0.5
(h) Student-0.6
(i) Student-0.7
Figure 13: t-SNE visualization results on the CMU-MOSI test set with the increasing data missing rates. We use red and blue to represent positive and negtive respectively.
Appendix B Confusion Matrices
We provide detailed confusion matrix results for the four-class and six-class IEMOCAP tasks. Similar quantities and data distributions at different rates demonstrate the model’s stability and robustness when dealing with incomplete data.
(a) Student-0.0
(b) Student-0.1
(c) Student-0.2
(d) Student-0.3
(e) Student-0.4
(f) Student-0.5
(g) Student-0.6
(h) Student-0.7
Figure 14: Visualization results of confusion matrices on the four-class IEMOCAP with the increasing data missing rates.
(a) Student-0.0
(b) Student-0.1
(c) Student-0.2
(d) Student-0.3
(e) Student-0.4
(f) Student-0.5
(g) Student-0.6
(h) Student-0.7
Figure 15: Visualization results of confusion matrices on the six-class IEMOCAP with the increasing data missing rates.
Appendix C Search Space
We select four computationally lightweight operations for our model:
  • 1.
    AttentionMixer:
    $MixMap=\frac{(X_{m}W_{q})(X_{m}W_{k})^{T}}{\sqrt{C}}$ (16)
    $output=(X_{m}W_{v})\otimes\mathrm{softmax}(MixMap)$ (17)
    $W_{q}$, $W_{k}$ and $W_{v}$ are trainable parameters.
  • 2.
    MLPMixer:
    $output=MLP(X_{m}^{T})^{T}$ (18)
  • 3.
    AvgPoolMixer:
    $output=\mathrm{AvgPool}(X_{m}^{T})^{T}$ (19)
  • 4.
    MaxPoolMixer:
    $output=\mathrm{MaxPool}(X_{m}^{T})^{T}$ (20)